# RIGNO: A Graph-based framework for robust and accurate operator learning for PDEs on arbitrary domains

### v1

@@ D.2.2 Wave-C-Sines
This dataset pertains to time-dependent solutions of the two-dimensional wave equation over a circular domain of radius $1$, centered at (0.5, 0.5)0.50.5(0.5,\,0.5)( 0.5 , 0.5 ). We set $c=2.0$, constant in time and in space, and impose homogeneous Dirichlet boundary conditions ($u=0$) on the boundary of the circle. The initial condition for $u$ at $t=0$ is specified as a superposition of sinusoidal modes:
@@
where $K=10$, $r=0.5$, and $a_{ij}\sim\mathcal{U}_{[-1,1]}$. We employ a Delaunay-based triangular mesh with characteristic length $0.015$, resulting in roughly $16{,}431$ nodes and $32{,}441$ elements. A time step of $\Delta t=0.001$ is used, and solutions are computed for $100$ time steps via a finite element solver. We then downsample the computed solution fields in time, keeping $21$ (including the initial condition) equally spaced time frames, which constitute the dataset.
@@ D.2.5 NS-SL
$v^{0}_{x}(x,y)$ =$=\begin{cases}\tanh\left(2\pi\frac{y-0.25}{\rho}\right)&\mbox{ for }y+\sigma_{\delta}(x)\leq\frac{1}{2}\\ \tanh\left(2\pi\frac{0.75-y}{\rho}\right)&\mbox{ otherwise}\end{cases},$ (24)
@@ D.2.6 NS-SVS
$\Gamma$ =$=\{(x,y)\in\mathbb{T}^{2}\mid y=\frac{1}{2}+0.2\sin(2\pi x)+\sum_{i=1}^{10}\alpha_{i}\sin(2\pi(x+\beta_{i}))\}.$ (30)
@@ D.2.10 Wave-Layer
In this dataset, the transonic flow is considered over different airfoils. The steady-state form of the compressible Euler equations (15) govern this system. The viscous effects are ignored and the far-field boundary conditions of $\rho_{\infty}=1$ , $p_{\infty}=1.0$, $M_{\infty}=0.8$ and zero angle of attack are considered. At the airfoil boundary, no-penetration conditions are considered. We refer the reader to [21] for more details about how the airfoil geometries are drawn. The flow is computed on a C-grid mesh with 200$\times$50 quadrilateral elements by a second-order implicit finite volume solver. The original dataset is enhanced by adding a distance field from the boundary of the airfoil which is used as the input of the model. The target value used in this work is the density $\rho$.
@@
D.2.12 Elasticity
@@ Appendix E Details of the Experiments
The hyperparameters of RIGNO-18 and RIGNO-12 are listed in Table E.1. The only difference between the two architectures is in the number of message passing blocks. The presented settings in this section are used in all the presented experiments, unless otherwise stated. A JAX [3] implementation of RIGNO is available on Github (https://github.com/camlab-ethz/rigno).
@@
4 4 Number of frequencies for encoding coordinates, i.e., $K_{\rm freq}$ in Equation (4).
@@
We train our models with 1024 training samples for all datasets; except for the AF dataset where 2048 samples are used in the trainings. We use a different set of 128 samples for validation, and 256 samples for testing. The test errors reported in the second part of Table 1 (uniform grid datasets) are exceptionally calculated on 240 test samples. We train our models for 2500 epochs on the time-dependent datasets, and for 3000 epochs on the time-independent datasets. We calculate the errors on the validation dataset regularly and keep the model parameters that give the lowest validation error. We use batch size 16, the AdamW [26] optimizer with weight decay $10^{-8}$, the mean squared error loss function, and cosine learning rate decay with linear warm-up period. The learning rate is warmed up in the first 2% training epochs from $10^{-5}$ to $2\times 10^{-4}$, and decays with cosine schedule to $10^{-5}$ in the next 88% of the epochs. In the last 10% epochs, it exponentially decays to $10^{-6}$. For all the time-dependent datasets, the time-derivative time stepping scheme is used due to its better performance (see Figure 3(c)). We use the all2all training strategy [12] with a warm-up period. We first train the model with all the pairs with lead time $2\Delta t$ for a few epochs, then add the pairs with lead time $4\Delta t$ and train for more epochs, and continue until all lead times are covered. The warm-up period for the all2all training takes 20% of the training epochs.
@@
Appendix F Baseline architectures
F.1 CNO

### v2

@@ D.2.2 Wave-C-Sines
This dataset pertains to time-dependent solutions of the two-dimensional wave equation over a circular domain of radius $1$, centered at (0.5, 0.5)(0.5,\,0.5). We set $c=2.0$, constant in time and in space, and impose homogeneous Dirichlet boundary conditions ($u=0$) on the boundary of the circle. The initial condition for $u$ at $t=0$ is specified as a superposition of sinusoidal modes:
@@
where $K=10$, $r=0.5$, and $a_{ij}\sim\mathcal{U}_{[-1,1]}$. We employ a Delaunay-based triangular mesh with characteristic length $0.015$, resulting in roughly 16,43116{,}431 nodes and 32,44132{,}441 elements. A time step of $\Delta t=0.001$ is used, and solutions are computed for $100$ time steps via a finite element solver. We then downsample the computed solution fields in time, keeping $21$ (including the initial condition) equally spaced time frames, which constitute the dataset.
@@ D.2.5 NS-SL
[…]
@@ D.2.6 NS-SVS
$\Gamma$ $=\{(x,y)\in\mathbb{T}^{2}\mid y=\frac{1}{2}+0.2\sin(2\pi x)+\sum_{i=1}^{10}\alpha_{i}\sin(2\pi(x+\beta_{i}))\}.$ (28)
@@ D.2.10 Wave-Layer
[…]
@@
D.2.12 Poisson-Gauss
Poisson’s equation (16) is considered in a unit square with homogeneous Dirichlet boundary conditions. the source term $f$ is composed of a superposition of $M$ random Gaussians as
$f(x,y)=\sum_{i=1}^{M}\exp(-\frac{(x-\mu_{x,i})^{2}+(y-\mu_{y,i})^{2}}{2\sigma_{i}^{2}}),$ (32)
where $M$ is a random integer drawn from a geometric distribution, $\mu_{x,i},\mu_{y,i}\sim\mathcal{U}_{[0,1]}$, and $\sigma_{i}\sim\mathcal{U}_{[0.025,0.1]}$. A few input-output pairs of this dataset are visualized in Figure D.13.
Figure D.13: Inputs (top row, source function) and outputs (bottom row, solution) of the Poisson-Gauss dataset. Each column corresponds to one sample in the dataset.
D.2.13 Elasticity
@@ Appendix E Details of the Experiments
A central design principle of RIGNO is to minimize the need for extensive hyperparameter tuning by employing a fixed set of hyperparameters across all tasks. To this end, we identified a robust configuration by tuning on a single internal dataset and subsequently applied the same hyperparameters to all other datasets without further adjustment. The size and performance of RIGNO are primarily influenced by only two key architectural hyperparameters: the number of processor blocks and the latent dimensionality of edge and node features. A comprehensive list of the hyperparameters used in RIGNO-18 and RIGNO-12 is provided in Table E.1. Since most of these hyperparameters require no task-specific tuning, the same configuration can be readily adopted for new problems or datasets. The only distinction between the two architectures lies in the number of message-passing blocks. Unless otherwise stated, the settings described in this section are used for all experiments. As shown in Table 1, both RIGNO-18 and RIGNO-12 achieve consistently strong performance across diverse benchmarks without the need for task-specific tuning.
@@
4 4 Number of frequencies for encoding coordinates, i.e., $K_{\rm freq}$.
@@
We train our models with 1024 training samples for all datasets; except for the AF dataset where 2048 samples are used in the trainings. We use a different set of 128 samples for validation, and 256 samples for testing. The test errors reported Table 2 are exceptionally calculated on 240 test samples. We train our models for 2500 epochs on the time-dependent datasets, and for 3000 epochs on the time-independent datasets. We calculate the errors on the validation dataset regularly and keep the model parameters that give the lowest validation error. We use batch size 16, the AdamW [27] optimizer with weight decay $10^{-8}$, the mean squared error loss function, and cosine learning rate decay with linear warm-up period. The learning rate is warmed up in the first 2% training epochs from $10^{-5}$ to $2\times 10^{-4}$, and decays with cosine schedule to $10^{-5}$ in the next 88% of the epochs. In the last 10% epochs, it exponentially decays to $10^{-6}$. For all the time-dependent datasets, the time-derivative time stepping scheme is used due to its better performance (see Figure 3(c)). We use the all2all training strategy [13] with a warm-up period. We first train the model with all the pairs with lead time $2\Delta t$ for a few epochs, then add the pairs with lead time $4\Delta t$ and train for more epochs, and continue until all lead times are covered. The warm-up period for the all2all training takes 20% of the training epochs.
Most of the experiments with RIGNO have been done using 2 NVIDIA GeForce RTX 3090 GPUs (24GB). Some experiments have been carried out with more than 2 (up to 8) GPUs in order to speed up the trainings. We estimate a total of 9’000 GPU hours for reproducing the presented experiments with RIGNO (including SM).
@@
Appendix F Further Experimental Results
There are many hyperparameters that can affect the training stability as well as different aspects of the performance of a trained RIGNO. In the experiments presented in this work, we fix a set of architectural choices that perform fairly well for all the datasets and training sizes. These hyperparameters are listed in Table E.1. Note that these hyperparameters do not correspond to the best-performing architecture, but rather they guarantee a reasonable performance with all the desired properties. In Section F.3, we show that the performance can be improved by using larger models. In the experiments of this chapter, we perturb one or a few of these hyperparameters while leaving the rest fixed in order to study how the performance of the model changes with respect to the changed hyperparameters. The trainings are carried out using the settings described in Section E, unless otherwise stated.
F.1 Training randomness
In order to quantify the dependence of the final results on training randomness, we train RIGNO five times, each time with a different seed for the pseudo-random number generator. We store the model with the lowest validation error during training as in all experiments, and check for fluctuations in the test error across all five repetitions of each experiment. In these experiments, we use the unstructured version of the datasets (random point clouds), we train using 512 training samples with batch size 8, and do not fine tune with fractional pairing. The results are summarized in Table F.1, where the maximum and minimum test errors across the repetitions are reported for each dataset. These fluctuations are attributed to the randomness in processes such as weight initializations, dataset shuffling, edge masking, and random sub-sampling of the coordinates. In the rest of the experiments, we only train each model once with a fixed training random seed, keeping in mind that the errors can slightly deviate because of training randomness.
Table F.1: Statistics of relative $L^{1}$ test errors with different random seeds. We train RIGNO-12 on 512 samples from each datasets. Each experiment is repeated 5 times.
Dataset Error [%]
Mean $\pm$ Standard deviation
NS-PwC 2.50 $\pm$ 0.157
NS-SVS 1.05 $\pm$ 0.040
F.2 Number of regional nodes
The number of regional nodes is independent from the given physical nodes and can be chosen arbitrarily. With more regional nodes, the support sub-regions will be smaller on average, leading to fewer edges in the encoder and the decoder. However, the processor steps become significantly more expensive. In order to investigate the effect of different numbers of regional nodes on the performance, we train a slightly modified RIGNO-18 on the NS-SL and the NS-SVS datasets. The test results are presented in Table F.2. With only 1024 regional nodes, the model already achieves a reasonable error for both datasets. With more regional nodes, the error follows a decreasing trend with both datasets, although less significant with NS-SVS. However, the inference times show that fewer regional nodes are favorable in terms of computational efficiency. RIGNO-18 is 5 times more time-efficient with 1024 regional nodes as compared to 16384 regional nodes. Moreover, the decreasing trend of the error breaks with too many regional nodes. With sub-sample factor 1.0 (16384 regional nodes), the support sub-regions become so small that most of them only contain one physical node and hence cannot reflect a good representation of the whole spatial region. In the main experiments of the current work, we have chosen a sub-sample factor of 4.0 to maintain a balance between accuracy and efficiency.
Table F.2: Test errors and inference times with different number of regional nodes. The experiments are conducted with a variation of RIGNO-18, trained on 256 training trajectories for 2000 epochs without all2all pairing. Although the sub-sampling factor for obtaining the regional nodes changes, multi-scale edges are always obtained based on sub-meshes of factor 2.0 until the regional mesh saturates.
Regional nodes Median relative $L^{1}$ error [%] Inference time [ms]
(sub-sample factor) NS-SL NS-SVS
1024 (16.0) 4.32 2.90 7.0
2048 (8.0) 3.96 2.93 9.0
4096 (4.0) 3.49 2.84 12.9
8192 (2.0) 3.33 2.83 28.3
16384 (1.0) 4.12 3.05 35.8
F.3 Scaling with model size
In order to evaluate how the performance of RIGNO scales with model size, we start with RIGNO-12 as the base architecture and change the latent sizes (all at the same time), the number of message passing steps in the processor, the overlap factors (both at the same time), and the number of times that the regional mesh is sub-sampled when processor edges are built (number of regional edge levels). We train these models on 256 samples of the NS-PwC dataset without fractional pairing and calculate the error at $t_{14}$ on 128 test samples by using the AR-2 rollout scheme (defined in Section C.2). The results are plotted in Figure F.1.
We observe that the latent sizes and number of processor steps have the largest effect on the model capacity. The error consistently falls down when these hyperparameters are scaled, and we expect to get even lower errors with larger RIGNO architectures. With larger overlap factors, although the number of edges (and hence the computational cost) increases rapidly, the rate of convergence is relatively low. Therefore, the authors recommend using the minimum overlap, especially for the encoder. With more regional mesh levels, on the other hand, the number of edges only increases logarithmically, which is also reflected in the inference times. With 4 regional mesh levels, we obtain 3.75% error, which is extremely lower than the 17.10% error with only a single regional mesh level. The inference time, however, is only increased by 14%. These results show the effectiveness of longer edges in the construction of the regional mesh. However, depending on the space resolution of the inputs, there is saturation limit for this hyperparameter. If we consider a downsampling factor of 4, with 16’000 input coordinates, for instance, we end up with a mesh with only 4 nodes after 6 downsampling steps. With so few nodes we can only get 6 additional edges at most, which explains why the decreasing trend in the error stops at 4 downsampling steps.
Figure F.1: Test errors at $t=t_{14}$ with different strategies for scaling model size. The experiments are carried out on the NS-PwC dataset.
F.4 Spatial continuity: resolution invariance
In Figure 3(a), we presented results of zero-shot super-resolution and zero-shot sub-resolution inference of RIGNO. The results show that RIGNO maintains its accuracy when inferred with higher resolutions but the accuracy drops with lower resolutions. When trained with edge masking, zero-shot generalization to lower resolutions is considerably improved. Generalization to higher resolutions is often a desirable feature since the model can be trained with lower resolutions with considerably less training costs, and be deployed for high-resolution inputs and outputs. In Figure F.2, we present the results of the resolution invariance test on MeshGraphNet. The results show that MeshGraphNet admits great dependence on the training resolution, with very large errors on any resolution except for the training resolution. This dependence is alleviated when MeshGraphNet is trained with edge masking. The accuracy is much less affected both with lower and higher resolutions. However, edge masking alone is not enough to make MeshGraphNet resolution invariant. With only 4 times higher resolutions, the error is almost doubled ($55.16\%$) as compared to the error with the training resolution ($30.36\%$).
Figure F.2: Resolution invariance test if MeshGraphNet with $64^{2}$ training resolution on the NS-PwC dataset. EM stands for edge masking. The y-axis shows the relative test error (at $t=t_{14}$) when the spatial resolution of the inputs and outputs varies from the training resolution.
F.5 Temporal continuity: generalization to unseen time steps
In order to assess the generalization of the trained models to unseen lead times $\tau$, we train RIGNO on the NS-PwC dataset with $2\Delta t$ time resolution, and test it on the test dataset with $\Delta t$ time resolution. With the incorporated mechanisms, namely time-conditioned normalization layers, all2all training, and fractional pairing fine-tuning, we expect to see a monotone increase in the error as we increase the lead time; which is very well captured in our tests as visualized in Figure F.3. Here, the lead time $\tau$ is fixed at each line in the plot, and the input time snapshot $t$ is changed. This explains why the first entries are missing for larger lead times. In the uppermost line ($\tau=9\Delta t$), for instance, the first entry corresponds to $t=0$, $\tau=9\Delta t$; the second entry to $t=1\Delta t$, $\tau=10\Delta t$; and so on.
In Figure F.3, the errors increase with larger lead times, which is expected since estimating the solution of the PDE with a higher lead time is inherently more difficult. In these experiments, we have removed the pairs with $\tau>8\Delta t$ from the all2all training pairs. Hence, $\tau=9\Delta t$ is considered as extrapolation in lead time. Since lead times $\tau=3\Delta t$, $\tau=5\Delta t$, and $\tau=7\Delta t$ are also not present in the all2all pairs, they are considered as interpolation in lead time. We can see that the monotonicity of the increase in the error is preserved for the extrapolated as well as the interpolated lead times. Moreover, the error increases in a controlled manner. The authors believe that this property is attributed to incorporation of time-conditioned normalization layers in the architecture and the use of time-derivative time-stepping strategy. However, with the smallest lead time $\tau=\Delta t$, it was observed that the monotonicity was not preserved without fine-tuning with the fractional pairing strategy (see Figure 3(b)).
Figure F.3: Direct test errors with different $\tau$ for the NS-PwC dataset. The vertical line indicates $t=t_{14}$, which is the last time snapshot seen by the model during training. Different lines denote different lead times. The gray lines correspond to lead times present in the training dataset; the green lines highlight interpolation in lead time; and the red lines indicate extrapolation in lead time. The model is trained on 512 samples with batch size 4.
F.6 Model uncertainty
The edge masking technique randomly disables a subset of the edges in each message-passing block. By adding randomness to the model, edge masking allows inferring the same model multiple times with the same input in order to generate an ensemble of estimates at a later time. The standard deviation of the generated ensemble can be interpreted as an indicator of model uncertainty on each output coordinate.
Figure F.4 visualizes the mean and the standard deviation of an ensemble of estimates at $t=t_{14}$ generated from input at $t=t_{0}$, along with the absolute error of the mean with respect to the ground-truth solution. More such visualizations are available in Section H. These visualizations show how the regions with the highest errors largely overlap with the regions where the ensemble has the highest deviations; which supports the interpretation of the deviations as a measure for model uncertainty, and shows that they are good indicators for inaccuracies in the estimates generated by the model.
Figure F.4: Mean and standard deviation of an ensemble of autoregressive estimates at $t_{14}$ for a test sample of the NS-SL dataset.
F.7 Robustness to noise
(a) Heat-L-Sines
(b) NS-PwC
Figure F.5: Autoregressive test errors introduced by the noise in the input (relative). The y-axes show the difference in the relative error with noisy and clean inputs. The solid lines (best) show the results with edge masking and overlap factors 2.0. The dashed lines show the results with the same overlap factors but without edge masking during training. The dotted lines show the results with edge masking and overlap factors 1.0. All models are trained on 256 trajectories, without all2all training, and without fractional pairing fine-tuning.
For each input in the test dataset, we compute its standard deviation $\sigma_{\rm input}$, and add a Gaussian noise with standard deviation $\sigma_{\rm noise}=C_{noise}\sigma_{\rm input}$, with a noise level $C_{noise}\in\{0.005,0.02\}$. Starting with a noisy input, we apply the operators autoregressively with $\tau_{\rm max}=2\Delta t$. The errors in the resulting trajectories are partially due to the input noise, and partially to the approximation error of the operator. In order to distinguish between these two, we also unroll the operator with clean input ($C_{noise}=0$), and subtract the resulting errors from the errors with noisy inputs. This procedure gives a measure for the part of the error that is introduced by the noise in the input, denoted by noise-induced error. We present the relative noise induced error with respect to the approximation error: $|{e_{\rm noisy}-e_{\rm clean}}|/{e_{\rm clean}}$, where $e$ represents the autoregressive test error at $t=t_{14}$. The results are visualized in Figure F.5. With $2\%$ input noise at $t_{0}$, the noise is very well dampened to only $0.5\%$ at $t_{14}$ of the Heat-L-Sines dataset. With larger support sub-regions, the regional nodes aggregate data from a larger pool of physical nodes. Therefore, they are less sensitive to noisy inputs, and a stronger noise dampening is observed. Edge masking also plays an important role in noise dampening. When larger support sub-regions are used along with edge masking, $2\%$ input noise is dampened below $0.2\%$ at $t_{14}$.
F.8 Out-of-distribution generalization
A central challenge in operator learning, and in supervised machine learning more broadly, is achieving strong generalization to out-of-distribution (OOD) inputs and outputs. These models often experience a degradation in performance when evaluated on samples that deviate from the training distribution. While some degree of performance decay is inevitable, a robust framework should maintain reasonable accuracy under mild to moderate distribution shifts. To evaluate RIGNO’s robustness in this regard, we train it on three datasets taken from [38], namely Smooth Transport, Darcy Flow, and Compressible Euler Equations. The trained models are then tested on 128 in-distribution (ID) and 128 OOD samples. The OOD samples differ from the training distribution in the location of perturbations, characteristic length scales, or geometry complexity. All models are trained for at least 2000 epochs using a batch size of 4 to ensure convergence, and all evaluation metrics are computed on unnormalized (raw) test data.
The results are summarized in Table F.3. Interestingly, RIGNO exhibits strong robustness to OOD samples in the Darcy Flow and Compressible Euler problems, but experiences a pronounced performance drop for the Smooth Transport problem when evaluated on OOD inputs. In the case of the Compressible Euler problem, the OOD samples feature airfoils with more complex geometries, while in Darcy Flow, the OOD inputs differ primarily in their characteristic length scales. Both cases can be interpreted as involving local variations in the structure of the input and output functions, to which RIGNO generalizes well. In contrast, the Smooth Transport problem introduces a non-local shift, as the center of the Gaussian perturbation changes significantly in the OOD samples, leading to the observed performance degradation.
Table F.3: Median relative $L^{1}$ errors [%] on 128 in-distribution (ID) and 128 out-of-distribution (OOD) test samples. The errors are calculated on unnormalized samples.
Dataset Training samples ID OOD
Darcy Flow 256 2.57 3.18
Smooth Transport 512 0.72 6.50
Compressible Euler 512 0.51 0.62
F.9 Pre-training and transfer learning
PDE foundation models have been proposed [13, 11, 44, 30] as a solution to overcome the high data demands of neural operators. In this context, we have assessed the generalization capabilities of RIGNO to new input and output distributions. To this end, we pre-train RIGNO on 1024 trajectories (2000 epochs, without all2all pairing) of one dataset (e.g., NS-Gauss) and transfer it to a different dataset (e.g., NS-PwC). We fine-tune the pre-trained RIGNO with only a few trajectories of the target dataset and compare the results with a randomly initialized RIGNO. We have chosen the NS-Gauss, NS-PwC, and NS-SL datasets, which have the same underlying PDE (incompressible Navier-Stokes equation) but correspond to very different data distributions (see Figures D.4, D.5, and D.6). The results are depicted in Figure F.6. For the NS-PwC dataset, we find that with only 4 trajectories, the transferred RIGNO pre-trained on NS-Gauss has an error of 13.1%, while a RIGNO trained from scratch requires 128 trajectories to achieve the same error, providing a 32x gain in sample efficiency with transfer learning. Notably, the gain is smaller (roughly 4x) when the model is pre-trained on the NS-SL dataset. Similarly, for the NS-SL dataset as the target dataset, transfer learning with 32 samples lead to an error of 4.7%, which is 7 times less than the trained-from-scratch RIGNO (33.4%) with 32 samples. These examples illustrate the ability of RIGNO to transfer very well out of distribution.
(a) Target dataset: NS-PwC.
(b) Target dataset: NS-SL.
Figure F.6: The effect of pre-training on data efficiency of RIGNO-12 with the unstructured NS datasets. All errors are relative test errors on 256 test samples.
F.10 Performance benchmarks
We benchmarked RIGNO against the best-performing baseline, GINO, in terms of training time, accuracy, inference time, and peak memory usage during inference. For training, we used the maximum feasible batch size for RIGNO and a fixed batch size of 64 for GINO. In cases involving varying geometries across training samples, such as in the Elasticity dataset, the original GINO implementation does not support batch processing, requiring a reduced batch size of 1. We evaluated two RIGNO-18 configurations with 96 and 128 latent features, respectively, and two GINO variants comprising 3 and 7 FNO layers.
Table F.4: Performance benchmarks on two unstructured datasets evaluated on a single NVIDIA GeForce RTX 3090 GPU (24GB). $T$: training time [s]; TE: test error [%]; IT: inference time [ms]; and PM: peak memory usage [MB].
Poisson-Gauss Elasticity
Architecture Size $T$ TE IT PM $T$ TE IT PM
GINO (3) 3.6M 1.33 5.24 7.8 370 18.2 3.36 4.1 80
GINO (7) 8.4M 1.72 8.01 10.1 406 27.1 3.94 6.4 116
RIGNO-18 (96) 1.4M 24.9 2.23 8.1 121 2.4 4.33 1.5 51
RIGNO-18 (128) 2.5M 28.4 2.09 12.7 159 3.1 3.71 1.9 64
The results, summarized in Table F.4, indicate that RIGNO trains more slowly than GINO on datasets with fixed geometries and input coordinates, but trains faster on datasets with variable geometries, primarily due to its support for batch processing. Although longer training times are a drawback, training constitutes a one-time cost, whereas inference speed and memory usage are typically more critical in practice. In this regard, RIGNO and GINO exhibit comparable inference times, yet RIGNO consistently achieves substantially higher accuracy. Furthermore, RIGNO is up to three times more memory-efficient than GINO. Owing to its multi-scale architecture and computational efficiency, RIGNO stands out as a powerful end-to-end GNN that delivers state-of-the-art accuracy and robustness, while remaining competitive with strong non-GNN baselines in both speed and memory consumption.
Appendix G Baseline Architectures
G.1 CNO